# SafeAuto: Knowledge-Enhanced Safe Autonomous Driving with Multimodal Foundation Models

### v1

@@ Abstract
Traditional autonomous driving systems often struggle to integrate high-level reasoning with low-level control, resulting in suboptimal and sometimes unsafe driving behaviors. The emergence of multimodal large language models (MLLMs), which can process both visual and textual data, presents an opportunity to unify perception and reasoning tasks within a single framework. However, effectively embedding precise safety knowledge into MLLMs for autonomous driving remains a significant challenge. To address this, we propose SafeAuto, a novel framework that enhances MLLM-based autonomous driving systems by incorporating both unstructured and structured knowledge. Specifically, we first introduce the Position-Dependent Cross-Entropy (PDCE) loss function, designed to improve the accuracy of low-level control signal predictions when numerical values are represented as text. Second, to ensure safe autonomous driving by explicitly integrating precise safety knowledge into the MLLM, we develop a reasoning component for SafeAuto. This component translates driving safety regulations into first-order logic rules (e.g., “red light stop”) and incorporates these rules into a probabilistic graphical model, such as a Markov Logic Network (MLN). The MLN is trained to verify the predicted next actions using environmental attributes identified by attribute recognition models (e.g., detecting a red light) to form the predicates. Additionally, we construct a Multimodal Retrieval-Augmented Generation (Multimodal RAG) model that leverages video data, control signals, and environmental attributes to learn more effectively from past similar driving experiences.
By integrating PDCE, MLN, and Multimodal RAG, SafeAuto significantly outperforms existing baselines across multiple datasets. This advancement paves the way for more accurate, reliable, and safer autonomous driving systems that effectively learn from experience, adhere to traffic regulations, and execute precise control actions.
@@ 1 Introduction
Firstly, for low-level action prediction, current approaches in adapting MLLMs generally follow two fashions. The first fashion treats the prediction of float numbers as text generation (Gruver et al., 2024; Xu et al., 2024), directly training the MLLM using cross-entropy (CE) loss for token prediction. Some variations (Brohan et al., 2023; Sima et al., 2023) of this method involve tokenizing the prediction range into several bins and adding new tokens for each bin into the LLM’s vocabulary, allowing the model to predict the corresponding bin token ID. However, these methods remain somewhat coarse compared to traditional regression techniques (Hu et al., 2023) using Mean Squared Error (MSE) loss. Alternatively, another fashion (Jin et al., 2024) employs a linear layer to decode the float number from the output hidden embeddings of the MLLM, enabling the use of MSE loss to train the model. While this approach may improve numerical accuracy, it compromises the autoregressive capability of the LLM, as the model can then only be purely used for numerical prediction and cannot perform any further QA-for example, handling high-level question-answering. Additionally, regarding high-level action prediction, a significant limitation of current methods is their inability to effectively utilize both structured and unstructured knowledge when making decisions. Specifically, existing approaches often focus solely on data-driven techniques, inadequately incorporating structured knowledge such as traffic rules and safety constraints. Although some methods (Sima et al., 2023; Mao et al., 2023; Wang et al., 2024) attempt to include traffic regulations by embedding them into the model’s context, this implicit approach is insufficient. Due to the inherent tendency of MLLMs to hallucinate, they may still generate unsafe or illegal actions. Meanwhile, while RAG (Lewis et al., 2020) has been employed in language models (Semnani et al., 2023; Zhang et al., 2024a) to mitigate issues like hallucination by incorporating relevant information from external sources, Yuan et al. (2024) first propose to combine the rich multimodal data inherent in autonomous driving contexts—such as videos, images, and control signals—to learn from past driving experiences as unstructured knowledge.
@@ 2 Related Work
Autonomous Driving with MLLM. The emergence of MLLMs enables unified end-to-end generation of both high-level and low-level outputs. Most of these works often treat numerical control signals as text, training models using token prediction with cross-entropy loss. For example, DriveGPT4 (Xu et al., 2024) just treats low-level control signals as text, fine-tuning an MLLM to sequentially predict high-level and low-level actions in a conversational manner using the BDD-X dataset. DriveLM-Agent (Sima et al., 2023), influenced by RT-2 (Brohan et al., 2023), discretizes waypoints into bins, expanding the tokenizer vocabulary accordingly and fine-tuning the BLIP-2 (Li et al., 2023). While this facilitates end-to-end training, it remains coarse compared to UniAD (Hu et al., 2023), which uses MSE loss. Time-LLM (Jin et al., 2024) decodes numerical predictions directly from output embeddings using a linear layer with MSE loss but diminishes the language model’s autoregressive capabilities, limiting high-level question-answering abilities. Additionally, Tan et al. (2024) suggest that employing the LLM backbone in this way does not enhance regression performance. In contrast, we propose a novel PDCE loss that adapts the cross-entropy loss for numerical training to behave more like MSE loss while preserving the model’s ability to perform high-level question-answering.
Safety Guarantee. Further advancements involve integrating perception and planning tools into the MLLM context. Agent-Driver (Mao et al., 2023) incorporates modules from UniAD into an MLLM framework, serving as a language agent for autonomous driving. OmniDrive (Wang et al., 2024) introduces a framework combining 3D perception, reasoning, and planning. However, these methods remain purely data-driven and lack explicit safety verification for generated actions. Given the safety-critical nature of autonomous driving, ensuring that output actions are safe and compliant with traffic rules is essential. To address this, we incorporate extracted knowledge—specifically structured traffic rules—into a probabilistic graphical model like a Markov Logic Network (MLN) for explicit post-safety verification, which has been widely used in previous work (Yang et al., 2022; Zhang et al., 2023). Besides, RAGDriver (Yuan et al., 2024) further enhances reasoning by retrieving similar driving experiences through triplet loss-based metric learning. We extend this approach by developing a more flexible and efficient retrieval system, directly training a joint embedding based on multimodal inputs to learn relative rankings from text similarity. Most importantly, we find that the incorporation of binary structured environmental predicates (e.g., the presence of a stop sign) from the previous reasoning components, namely MLNs, significantly improves retrieval performance.
@@ 4 Experiments
Baselines. (a) On the BDD-X dataset, we compare our method with several baselines: (1) ADAPT (Jin et al., 2023), a state-of-the-art video transformer-based method that provides high-level and low-level answers using two separate branches; (2) DriveGPT4 (Xu et al., 2024), the first work to provide both high-level action descriptions and low-level vehicle control signals in an end-to-end fashion using an MLLM; and (3) RAGDriver (Yuan et al., 2024), a state-of-the-art method that leverages triplet loss to train multimodal retrieval models for autonomous driving. (b) For the DriveLM dataset, we use: (1) DriveLM-Agent, the current state-of-the-art method that employs graph-based visual question answering to improve high-level responses and uses motion tokenization for low-level prediction; (2) UniAD (Hu et al., 2023), the state-of-the-art method on the nuScenes dataset used here for comparing low-level predictions—-we consider two versions: UniAD (Full), which utilizes the entire historical video input, and UniAD (Single), a variant modified to use only the current frame’s input for a fair comparison; and (3) BLIP-RT-2, which fine-tunes BLIP-2 (Li et al., 2023) on the DriveLM data and utilizes trajectory tokenization as proposed in RT-2 (Brohan et al., 2023).
@@ 5 Ablation Study
Contribution from Each Module. The influence of each module on both high-level and low-level predictions is demonstrated in Section C.3. To accurately measure improvements in action performance, we introduce a new metric called high-level action predicate accuracy for the BDD-X dataset, which converts high-level action descriptions to one of 16 predefined actions using GPT4o strictly and measures accuracy. Our results reveal that: (1) PDCE loss markedly enhances low-level prediction while preserving high-level prediction performance; (2) post-safety verification via MLN still corrects certain unsafe actions, even though the base model is conservative; the impact of each module on the rates of rule violation is detailed in Section C.5. (3) the use of Multimodal RAG significantly enhances performance, increasing high-level action predicate accuracy by 30% on BBD-X, and 14% on DriveLM.
@@
Case Study on Post-Safety Verification with MLN. In the BDD-X dataset, the critical traffic rule and in the DriveLM dataset, the key rule is . These rules carry the highest weights in their respective MLNs. A case of MLN correcting aggressive driving behavior is illustrated in Figure 8.
@@ Acknowledgment
This work is partially supported by the National Science Foundation under grant No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA GARD, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, the Alfred P. Sloan Fellowship, the Meta research award, the AI Safety Fund, and the eBay research award.
@@ Appendix C Experiements on Ablation Study
| Method | K | Action | Justification | |||||
| B4 | C | M | Acc | B4 | C | M | ||
| Base | - | 30.8 | 221.5 | 29.2 | 61.75 | 7.8 | 85.4 | 13.2 |
| RAG w/o EP | 1 | 29.4 | 219.2 | 28.5 | 59.06 | 7.3 | 74.8 | 12.6 |
| RAG w/o EP | 2 | 29.7 | 218.6 | 28.7 | 59.91 | 7.3 | 73.7 | 12.5 |
| RAG w/ EP | 1 | 38.1 | 334.8 | 35.4 | 91.47 | 8.8 | 89.2 | 13.5 |
| RAG w/ EP | 2 | 38.2 | 334.8 | 35.3 | 91.00 | 9.4 | 95.5 | 13.9 |
C.1 Impact of varying
The impact of varying values on low-level predictions in the BDD-X dataset is demonstrated in Figure 4.
C.2 Impact of the for retrieval examples
The impact of varying for retrieval examples used on the BDD-X dataset is demonstrated in Table 10.
C.3 Contribution from different module
The contributions of different modules to the final high-level performance on the BDD-X dataset are shown in Table 12, and on the DriveLM dataset in Table 12. While Table 13 presents an ablation study evaluating the contribution of each module in SafeAuto to the low-level control signal prediction on the BDD-X dataset. Interestingly, we find that the MLN reasoning and RAG modules have only a minimal impact on the low-level prediction accuracy, with the primary improvement stemming from the PDCE loss, as expected. Additionally, we observe that incorporating RAG slightly increases the RMSE for speed prediction but decreases the RMSE for course prediction.
| Method | Action | Justification | |||||
| B4 | C | M | Acc | B4 | C | M | |
| Base | 30.8 | 221.5 | 29.2 | 61.75 | 7.8 | 85.4 | 13.2 |
| PDCE | 31.4 | 231.4 | 29.3 | 61.94 | 7.9 | 84.2 | 13.2 |
| PDCE + MLN | 31.5 | 232.2 | 29.4 | 62.97 | 7.9 | 84.5 | 13.2 |
| PDCE + RAG | 38.2 | 334.8 | 35.3 | 91.00 | 9.4 | 95.5 | 13.9 |
| PDCE + MLN + RAG | 38.6 | 337.4 | 35.5 | 92.18 | 9.4 | 96.0 | 14.0 |
| Method | High-Level Behavior | Motion | ||
| Acc | Speed | Steer | ADE | |
| Base | 60.58 | 64.67 | 80.29 | 0.86 |
| PDCE | 63.21 | 67.88 | 79.27 | 0.85 |
| PDCE + MLN | 66.86 | 71.39 | 80.29 | 0.85 |
| PDCE + RAG | 74.01 | 79.27 | 81.61 | 0.84 |
| PDCE + MLN + RAG | 74.60 | 79.85 | 81.90 | 0.84 |

### v2

@@ Abstract
Traditional autonomous driving systems often struggle to connect high-level reasoning with low-level control, leading to suboptimal and sometimes unsafe behaviors. Recent advances in multimodal large language models (MLLMs), which process both visual and textual data, offer an opportunity to unify perception and reasoning. However, effectively embedding precise safety knowledge into MLLMs for autonomous driving remains a significant challenge. To address this, we propose SafeAuto, a framework that enhances MLLM-based autonomous driving by incorporating both unstructured and structured knowledge. First, we introduce a Position-Dependent Cross-Entropy (PDCE) loss to improve low-level control signal predictions when values are represented as text. Second, to explicitly integrate safety knowledge, we develop a reasoning component that translates traffic rules into first-order logic (e.g., “red light stop”) and embeds them into a probabilistic graphical model (e.g., Markov Logic Network) to verify predicted actions using recognized environmental attributes. Additionally, our Multimodal Retrieval-Augmented Generation (RAG) model leverages video, control signals, and environmental attributes to learn from past driving experiences. Integrating PDCE, MLN, and Multimodal RAG, SafeAuto outperforms existing baselines across multiple datasets, enabling more accurate, reliable, and safer autonomous driving. The code is available at https://github.com/AI-secure/SafeAuto.
@@ 1 Introduction
Firstly, for low-level action prediction, current approaches in adapting MLLMs generally follow two fashions. The first fashion treats the prediction of float numbers as text generation (Gruver et al., 2024; Xu et al., 2024), directly training the MLLM using cross-entropy (CE) loss for token prediction. Some variations (Brohan et al., 2023; Sima et al., 2023) of this method involve tokenizing the prediction range into several bins and adding new tokens for each bin into the LLM’s vocabulary, allowing the model to predict the corresponding bin token ID. However, these methods remain somewhat coarse compared to traditional regression techniques (Hu et al., 2023) using Mean Squared Error (MSE) loss. Alternatively, another fashion (Jin et al., 2024b) employs a linear layer to decode the float number from the output hidden embeddings of the MLLM, enabling the use of MSE loss to train the model. While this approach may improve numerical accuracy, it compromises the autoregressive capability of the LLM, as the model can then only be purely used for numerical prediction and cannot perform any further QA-for example, handling high-level question-answering. Additionally, regarding high-level action prediction, a significant limitation of current methods is their inability to effectively utilize both structured and unstructured knowledge when making decisions. Specifically, existing approaches often focus solely on data-driven techniques, inadequately incorporating structured knowledge such as traffic rules and safety constraints. Although some methods (Sima et al., 2023; Mao et al., 2023; Wang et al., 2024) attempt to include traffic regulations by embedding them into the model’s context, this implicit approach is insufficient. Due to the inherent tendency of MLLMs to hallucinate, they may still generate unsafe or illegal actions. Meanwhile, while RAG (Lewis et al., 2020) has been employed in language models (Semnani et al., 2023; Zhang et al., 2024a) to mitigate issues like hallucination by incorporating relevant information from external sources, Yuan et al. (2024) first propose to combine the rich multimodal data inherent in autonomous driving contexts—such as videos, images, and control signals—to learn from past driving experiences as unstructured knowledge.
@@ 2 Related Work
Autonomous Driving with MLLM. The emergence of MLLMs enables unified end-to-end generation of both high-level and low-level outputs. Most of these works often treat numerical control signals as text, training models using token prediction with cross-entropy loss. For example, DriveGPT4 (Xu et al., 2024) just treats low-level control signals as text, fine-tuning an MLLM to sequentially predict high-level and low-level actions in a conversational manner using the BDD-X dataset. DriveLM-Agent (Sima et al., 2023), influenced by RT-2 (Brohan et al., 2023), discretizes waypoints into bins, expanding the tokenizer vocabulary accordingly and fine-tuning the BLIP-2 (Li et al., 2023). While this facilitates end-to-end training, it remains coarse compared to UniAD (Hu et al., 2023), which uses MSE loss. Time-LLM (Jin et al., 2024b) decodes numerical predictions directly from output embeddings using a linear layer with MSE loss but diminishes the language model’s autoregressive capabilities, limiting high-level question-answering abilities. Additionally, Tan et al. (2024) suggest that employing the LLM backbone in this way does not enhance regression performance. In contrast, we propose a novel PDCE loss that adapts the cross-entropy loss for numerical training to behave more like MSE loss while preserving the model’s ability to perform high-level question-answering.
Safety Guarantee. Providing safety guarantees (Li et al., 2022; Zhang et al., 2023a) is always a fundamental concern in machine learning, especially in safety-critical applications like autonomous driving. Further advancements involve integrating perception and planning tools into the MLLM context. Agent-Driver (Mao et al., 2023) incorporates modules from UniAD into an MLLM framework, serving as a language agent for autonomous driving. OmniDrive (Wang et al., 2024) introduces a framework combining 3D perception, reasoning, and planning. However, these methods remain purely data-driven and lack explicit safety verification for generated actions. Given the safety-critical nature of autonomous driving, ensuring that output actions are safe and compliant with traffic rules is essential. To address this, we incorporate extracted knowledge—specifically structured traffic rules—into a probabilistic graphical model like a Markov Logic Network (MLN) for explicit post-safety verification, which has been widely used in previous work (Yang et al., 2022; Zhang et al., 2023b). Besides, RAGDriver (Yuan et al., 2024) further enhances reasoning by retrieving similar driving experiences through triplet loss-based metric learning. We extend this approach by developing a more flexible and efficient retrieval system, directly training a joint embedding based on multimodal inputs to learn relative rankings from text similarity. Most importantly, we find that the incorporation of binary structured environmental predicates (e.g., the presence of a stop sign) from the previous reasoning components, namely MLNs, significantly improves retrieval performance.
@@ 4 Experiments
| Method | Action | Justification | |||||
| B4 | C | M | Acc | B4 | C | M | |
| Base | 30.8 | 221.5 | 29.2 | 61.75 | 7.8 | 85.4 | 13.2 |
| PDCE | 31.4 | 231.4 | 29.3 | 61.94 | 7.9 | 84.2 | 13.2 |
| PDCE + MLN | 31.5 | 232.2 | 29.4 | 62.97 | 7.9 | 84.5 | 13.2 |
| PDCE + RAG | 38.2 | 334.8 | 35.3 | 91.00 | 9.4 | 95.5 | 13.9 |
| PDCE + MLN + RAG | 38.6 | 337.4 | 35.5 | 92.18 | 9.4 | 96.0 | 14.0 |
| Method | High-Level Behavior | Motion | ||
| Acc | Speed | Steer | ADE | |
| Base | 60.58 | 64.67 | 80.29 | 0.86 |
| PDCE | 63.21 | 67.88 | 79.27 | 0.85 |
| PDCE + MLN | 66.86 | 71.39 | 80.29 | 0.85 |
| PDCE + RAG | 74.01 | 79.27 | 81.61 | 0.84 |
| PDCE + MLN + RAG | 74.60 | 79.85 | 81.90 | 0.84 |
@@
Baselines. (a) On the BDD-X dataset, we compare our method with several baselines: (1) ADAPT (Jin et al., 2023), a state-of-the-art video transformer-based method that provides high-level and low-level answers using two separate branches; (2) TimeLLM (Jin et al., 2024a), which repurposes frozen large language models for general time series forecasting by reprogramming numerical inputs into text-based patches; (3) DriveGPT4 (Xu et al., 2024), the first work to provide both high-level action descriptions and low-level vehicle control signals in an end-to-end fashion using an MLLM; and (4) RAGDriver (Yuan et al., 2024), a state-of-the-art method that leverages triplet loss to train multimodal retrieval models for autonomous driving. (b) For the DriveLM dataset, we use: (1) DriveLM-Agent, the current state-of-the-art method that employs graph-based visual question answering to improve high-level responses and uses motion tokenization for low-level prediction; (2) UniAD (Hu et al., 2023), the state-of-the-art method on the nuScenes dataset used here for comparing low-level predictions—-we consider two versions: UniAD (Full), which utilizes the entire historical video input, and UniAD (Single), a variant modified to use only the current frame’s input for a fair comparison; and (3) BLIP-RT-2, which fine-tunes BLIP-2 (Li et al., 2023) on the DriveLM data and utilizes trajectory tokenization as proposed in RT-2 (Brohan et al., 2023).
@@ 5 Ablation Study
| Method | K | Action | Justification | |||||
| B4 | C | M | Acc | B4 | C | M | ||
| Base | - | 30.8 | 221.5 | 29.2 | 61.75 | 7.8 | 85.4 | 13.2 |
| RAG w/o EP | 1 | 29.4 | 219.2 | 28.5 | 59.06 | 7.3 | 74.8 | 12.6 |
| RAG w/o EP | 2 | 29.7 | 218.6 | 28.7 | 59.91 | 7.3 | 73.7 | 12.5 |
| RAG w/ EP | 1 | 38.1 | 334.8 | 35.4 | 91.47 | 8.8 | 89.2 | 13.5 |
| RAG w/ EP | 2 | 38.2 | 334.8 | 35.3 | 91.00 | 9.4 | 95.5 | 13.9 |
Contribution from Each Module. The contributions of different modules to both high-level and low-level performance on the DriveLM dataset are shown in Table 5, while the high-level results for the BDD-X dataset are presented in Table 5. The contributions of each module in SafeAuto to low-level control signal prediction on the BDD-X dataset are deferred to Section C.1. To accurately measure improvements in action performance, we introduce a new metric called high-level action predicate accuracy for the BDD-X dataset, which converts high-level action descriptions to one of 16 predefined actions using GPT4o strictly and measures accuracy. Our results reveal that: (1) PDCE loss markedly enhances low-level prediction while preserving high-level prediction performance; (2) post-safety verification via MLN still corrects certain unsafe actions, even though the base model is conservative; the impact of each module on the rates of rule violation is detailed in Section C.3. (3) the use of Multimodal RAG significantly enhances performance, increasing high-level action predicate accuracy by 30% on BBD-X, and 14% on DriveLM.
@@
Case Study on Post-Safety Verification. In the BDD-X dataset, the critical traffic rule and in the DriveLM dataset, the key rule is . These rules carry the highest weights in their respective MLNs. A case of MLN correcting aggressive driving behavior is illustrated in Figure 8.
@@ Acknowledgment
This work is partially supported by the National Science Foundation under grant No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, and Schmidt Science.
@@ Appendix C Experiements on Ablation Study
C.1 Contribution from different module
Table 13 presents an ablation study evaluating the contribution of each module in SafeAuto to the low-level control signal prediction on the BDD-X dataset. Interestingly, we find that the MLN reasoning and RAG modules have only a minimal impact on the low-level prediction accuracy, with the primary improvement stemming from the PDCE loss, as expected. Additionally, we observe that incorporating RAG slightly increases the RMSE for speed prediction but decreases the RMSE for course prediction.